# A Measurement of Solar Neutrino Flux using PandaX-4T Electron Recoil Data Xiaoying Lu

### v1

@@
A Measurement of Solar Neutrino Flux using PandaX-4T Electron Recoil Data
@@
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Changbo Fu
Mengting Fu
School of Physics, Peking University, Beijing 100871, China
@@
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, Henan 450001, China
@@
Linhui Gu
@@
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Yanlin Huang
School of Medical Instrument and Food Engineering, University of Shanghai for Science and Technology, Shanghai 200093, China
@@
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
jianglai.liu@sjtu.edu.cnTsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Andi Tan
@@
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jingkai Xia
@@
Pengwei Xie
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
yanbinbin@sjtu.edu.cnTsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Yukun Yao
@@
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Qibin Zheng
School of Medical Instrument and Food Engineering, University of Shanghai for Science and Technology, Shanghai 200093, China
@@ Abstract
The proton-proton () fusion chain dominates the neutrino production from the Sun. The uncertainty of the predicted neutrino flux is at the sub-percent level, whereas that of the best measurement is . Furthermore, no direct measurement has been made via electron scattering below 165 keV recoil energy. Using the PandaX-4T commissioning data with
0.63 tonneyear of exposure, we perform the first measurement of solar neutrinos in the recoil energy range from 24 to 144 keV. The neutrino flux is measured to be , consistent with Standard Solar Model and existing measurements.
According to the Standard Solar Model (SSM) [1, 2, 3], approximately 99% of the solar power comes from a series of reactions fusing hydrogen into helium.
Neutrinos emitted in the initial fusion of two protons to a deuteron constitute roughly 91% of the total solar neutrino flux. These neutrinos are commonly referred to as the neutrinos.
The sub-dominant neutrinos from electron capture of Be account for an additional 7% of the flux.
The weakly interacting neutrinos carry original information from solar fusion reactions.
@@
The solar luminosity has been measured to a 0.6% precision.
In contrast, the neutrino flux has only been measured by Borexino with a precision of with recoil energy threshold at 165 keV [6, 7].
A number of multi-tonne-scale liquid xenon (LXe) experiments including PandaX-4T [8], LZ [9] and XENONnT [10] are under operation to search for dark matter particles, coherent scattering of solar neutrinos on xenon nuclei [11, 12], and possible abnormal magnetic moments of neutrinos [13, 14, 15] in the few or few-tens of keV-scale electron or nuclear recoil energy.
By design, these detectors also effectively cover over an electron recoil energy up to several hundred keV, spanning the majority of region of interest (ROI) for the neutrino flux.
Figure 1: Neutrino-electron elastic scattering spectrum of solar neutrinos with xenon electrons.
If the binding energy of atomic electrons (AE) is considered (red curve), the scattering event rate in our ROI is lower than the free electron (FE) scenario (green curve).
Adopted from Ref. [16].
@@
The Be neutrinos are mono-energetic at 384 keV (approximately 10%) and 862 keV (90%).
The expected event rate per unit recoil energy is
where is the number of target electrons, is the neutrino flux as a function of neutrino energy, () is the oscillation probabilities of into the flavor j, and is the differential cross-section.
Fig. 1 shows the energy spectrum of solar and Be neutrinos in a LXe detector, as given in Ref. [16].
When the xenon atomic effects are considered (adopted later in this paper), the rate in the ROI is suppressed by a few percent compared to the free electron scenario.
In this paper, we report the first measurement of the neutrino flux via their elastic scattering on electrons using the PandaX-4T detector located in the B2 hall of the China Jinping Underground Laboratory [17, 18].
The sensitive target of PandaX-4T has 3.69 tonne liquid xenon in a cylindrical dual-phase xenon time projection chamber (TPC) with 118.5 cm in diameter and 116.8 cm in height [19].
@@
This analysis selects different data from the existing analyses of the first commissioning data release [8, 11, 20].
The energy ROI is from 24 keV to 144 keV, with more than 60% of the solar neutrinos included.
The lower boundary is chosen to be above the dark matter particle search region and the upper boundary to avoid the prominent 163.9 keV peak from Xe produced in the neutron calibration runs.
@@
Therefore, we can recover approximately 9.5% exposure previously excluded in low energy analysis due to elevated detector noise.
On the contrary, 8.4 days of data following a neutron calibration run is removed because of a high concentration of activated Xe and I.
In the end, a total of 86.5 days is used for this analysis.
The PMT gains are calibrated with a newly implemented “rolling gain” approach for PMTs, which calculates PMT gains by fitting single photoelectron (SPE) spectra in individual PMTs run by run, adopted already in recent Ref. [19].
@@
To avoid biases in our data selections, cuts are first determined with LED-gain calibration data, validated with 16% of rolling gain physics data, and finally applied to the entire data set.
The quality cut variables are inherited from low energy analysis [8], but cut parameters are modified
to suit the different energy windows.
@@
The solar neutrino signals scattering on electrons are primarily single-site (SS) events within the ROI.
@@
The SS efficiency is , calculated using the Rn calibration data.
A consistent efficiency is obtained using simulated data.
@@
The reconstruction uniformity is confirmed with the diffusive Kr calibration source injected into the TPC [21].
The mono-energetic peak of Kr is also used to generate an energy response map, and the corresponding correction procedure follows Ref. [8].
@@
The FV selection is the same as in the low energy analysis geometrically, and a total of tonne of LXe in the center of the TPC is used.
@@
Low energy background events from Rn chain are mainly the decay of Pb to the ground state of Bi [19].
The event ratio, defined as the fraction of Pb events in the ROI per decay, is determined from a dedicated calibration run with Pb events accumulated in about 11 days.
@@
The number of Pb events in the ROI is .
The decay of Bi is a sub-dominate background component because it is rejected effectively by the Bi–Po coincidence analysis.
The Bi–Po pairs are identified by an event following a event in the subsequent 5 ms window.
In the time window, the number of randomly paired events is negligible.
We remove the identified Bi–Po pairs from our spectrum, and the contribution from Bi is no longer included in the spectrum fit.
Similar to Pb, the expected contribution from decay of Pb in Rn decay chain is determined from spectrum fit from higher energy analysis [19].
The maximum deviation of Pb/Po throughout the calibration period and stable decay period is considered as the systematic uncertainty.
The expected number of events in the ROI is .
The concentration of Kr is 0.52 0.27 parts per trillion [19].
The expected contribution of Kr is 489 254 in the ROI.
The levels of radioactivity from PMTs and detector vessels have been determined from the wide energy spectrum fit in Ref. [20].
The expected contribution in ROI is events.
The expected background from Xe is in the ROI, given by the half-life measured by PandaX-4T experiment with an exposure of 15.5 kgyear of Xe isotope [20].
The contributions of Xe, Xe, I, and Xe are free and fitted in the final spectral analysis, as there are no constraints in our data outside the ROI.
Xe from neutron activation during calibration contributes to the ROI with a distinctive rising slope starting at 81 keV in the energy spectrum.
Cosmogenic Xe, introduced when some xenon exposed at the Earth’s surface was filled into the detector during the data taking, contributes to the background with K-shell electron capture around 33.2 keV.
Neutron-activated Xe decays quickly to
the relatively long-lived I with a half-life of 59.4 days.
I electron capture results in energy release of 67.3 keV (80% K-shell), 40.4 keV (16% L-shell), and 36.5 keV (3.5% M-shell) in the ROI.
Thanks to the efficient removal capability of the xenon purification and circulation system, the effective lifetime of I in the TPC is significantly reduced and cannot be constrained by the isotopic lifetime.
Two-neutrino double-electron capture of Xe (abundance = 0.095%, Q = 2857 keV) deposits energy within our ROI at 64.3 keV and 32.3 - 37.3 keV [25, 26].
Table 1: Expected and fitted background and signal counts in the ROI.
@@
The contributions from Pb, Pb, Kr, detector material, and Xe are constrained using the uncertainties of the expected counts as described earlier and listed in Table 1.
Multiple mono-energetic peaks around 67 keV (33 keV) are grouped and modeled using a Gaussian function.
The means of the Gaussian functions are free in the fit, while the sigmas are constrained by a fractional uncertainty of 20% of the measured resolutions in Fig. 2.
The contributions of Xe and two Gaussian peaks are free-floating in the fit.
The sum of solar and Be neutrino signals are combined into a single fit component.
The spectrum fit result is shown in Fig. 4.
The number of solar +Be neutrino signals is 297 260.
@@
The background contributions from the material, Kr, Pb, and Pb are consistent with the expected values.
The fitted uncertainty includes a statistical () and a systematic () uncertainty of 107 and 237 events, respectively, extracted by comparing the signal uncertainties with and without the penalty terms for the systematic constraints.
A similar ad hoc approach is used to extract the contribution from individual systematic effects later, assuming that extracted uncertainty is uncorrelated with the rest.
Xe (free) is the largest background source, however its contribution to is relatively insignificant at 43 events due to its characteristic spectral shape.
Table 2: Summary of contributions to systematic uncertainties. The contribution of each constrained parameter is extracted by turning on/off the corresponding penalty term in the fitter, assuming no correlation with other constraints. As a result, the quadrature combination of those rows is slightly different from the subtotal obtained directly by the fitter with all correlations properly built in.
@@
32
Pb spectrum
82
Pb spectrum
17
Kr spectrum
6
@@
277
Table 2 lists all the components of the systematic uncertainties.
Among the constrained backgrounds, Kr, Pb and Pb dominate the contribution to with an uncertainty of 203, 87, and 69 events, respectively.
The constraints of two Gaussian peak widths give an uncertainty of 37 and 33 events, respectively.
The uncertainties of data selection, including FV, quality cuts, and SS fraction, affect the uncertainty of the signal number proportionally with 30 events.
Additional sources of systematic uncertainties are evaluated manually ().
The uncertainty from the reconstructed energy scale is introduced by shifting the energy spectrum by keV and comparing the new fit results with the baseline fit, with a result of 111 events. To evaluate the contribution due to the energy resolution uncertainty,
we perform the fit using PDFs with different energy resolution values.
A difference in energy resolution, where is from the aforementioned resolution function, propagates into an uncertainty of 19 events.
The fit range uncertainty is evaluated to be 32 events by varying the ROI to 25 – 143 keV and 23 – 145 keV.
The uncertainty introduced by Pb spectrum shape is calculated by changing the theoretical spectrum in the baseline fit to the shape measured in calibration data, leading to a result of 82 events.
Similarly, the spectral shape differences between Geant4 [28] and a recent theoretical calculation of Pb and Kr [24] are also considered, and the introduced systematic uncertainties are found to be 17 and 6 respectively.
In total, we get an overall systematic uncertainty of 277 events with 144 from .
Figure 4: (Color online) Result of the spectrum fit and the corresponding residual plot in the bottom panel. The solar +Be neutrino is shown as the red line.
@@
The (number of degrees of freedom, defined as the number of bins minus the number of free fit parameters) is shown in the legend.
The final number of +Be neutrinos is from our analysis.
The corresponding solar neutrino flux is , taking into account of the expected ratio of the and Be fluxes from SSM and their contributions in the ROI.
@@
This analysis represents the first measurement for solar neutrino at low recoil energy and demonstrates the potential of multi-tonne-scale LXe detectors for solar neutrino studies in a completely new energy window.
@@
Additional efforts, such as replacing TPC materials and circulation pumps, are being implemented for background control.
Additional Rn calibration data and improved detector response with upgraded PMT readout circuit boards are expected to significantly reduce systematic uncertainties.
Assuming radon and krypton levels of 3.5 Bq/kg and 0.25 ppt, respectively, negligible activation of xenon isotopes, and imposing a 5% constraint on all backgrounds, PandaX-4T can measure the solar neutrinos with an uncertainty of 30% with 6 tonneyear exposure.
Figure 5: The red dot and line show the result of solar neutrino flux, compared with the measurement from Borexino [7] and PandaX-4T projected measurement with 6 tonneyear exposure.
@@
This project is supported in part by the grants from National Science Foundation of China (Nos. 12090060, 12090063, 12105052, 12005131, 11905128, 11925502), and by the Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100).
@@
Bellini et al. [2014]G. Bellini et al. (BOREXINO), Neutrinos from the primary
proton–proton fusion process in the Sun, Nature 512, 383 (2014).
@@
Yan et al. [2023]X. Yan et al. (PandaX), Searching for Two-Neutrino and Neutrinoless Double Beta
Decay of Xe with the PandaX-4T Experiment (2023), arXiv:2312.15632
[nucl-ex] .

### v2

# Measurement of Solar Neutrino Flux using Electron Recoil Data from PandaX-4T Commissioning Run Xiaoying Lu

@@
Measurement of Solar Neutrino Flux using Electron Recoil Data from PandaX-4T Commissioning Run
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yunhua Chen
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
jianglai.liu@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Manbin Shen
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Xu Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuan Wu
@@
Kaizhi Xiong
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
yanbinbin@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
@@
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
@@ Abstract
The proton-proton () fusion chain dominates the neutrino production from the Sun.
The uncertainty of the predicted neutrino flux is at the sub-percent level, whereas that of the best measurement is .
In this paper, we present the first result to measure the solar neutrinos in the electron recoil energy range from 24 to 144 keV, using the PandaX-4T commissioning data with 0.63 tonneyear exposure.
The neutrino flux is determined to be , consistent with Standard Solar Model and existing measurements, corresponding to a flux upper limit of at 90% C.L..
According to the Standard Solar Model (SSM) [1, 2, 3], approximately 99% of solar power comes from a series of reactions fusing hydrogen into helium.
Neutrinos emitted in the initial fusion of two protons to a deuteron constitute roughly 91% of the total solar neutrino flux. These neutrinos are commonly referred to as the neutrinos, with a theoretical flux uncertainty at the sub-percent level.
The sub-dominant neutrinos from electron capture of 7Be account for an additional 7% of the flux.
@@
The first observation of neutrino flux is achieved using radiochemical material of gallium by GALLEX/GNO [6] and SAGE [7]. The first real-time detection is made by Borexino with now a state-of-the-art precision of and an electron recoil energy threshold of 165 keV [8, 9].
In this paper, we present a new result to measure neutrino flux in real-time by PandaX-4T within a previously unaccessible recoil energy window between 24 to 144 keV. It also represents the first such measurement using a detector primarily designed for dark matter direct detection.
A number of multi-tonne-scale liquid xenon (LXe) experiments, including PandaX-4T [10], LZ [11], and XENONnT [12] are under operation to search for dark matter particles, coherent scattering of solar neutrinos on xenon nuclei [13, 14], and possible abnormal magnetic moments of neutrinos [15, 16, 17] in the few or few-tens of keV-scale electron or nuclear recoil energy.
By design, these detectors also effectively cover an electron recoil energy up to several hundred keV, spanning over most of the energy region of the neutrino.
Figure 1: Neutrino-electron elastic scattering spectrum of solar neutrinos with liquid xenon.
If the binding energy of atomic electrons (AE) is considered [18] (red curve), the scattering event rate in our region of interest is lower than the free electron (FE) scenario (green curve). The more complete treatment using the relativistic random phase approximation (RRPA) is shown in cyan, but only for a recoil energy less than 30 keV.
The 7Be neutrinos are mono-energetic at 384 keV (approximately 10%) and 862 keV (90%).
@@
The expected electron recoil event rate per unit recoil energy is
where is the number of target electrons, is the neutrino flux as a function of neutrino energy, () is the oscillation probabilities of into flavor j and is the differential cross-section.
Fig. 1 shows the electron recoil energy spectrum of solar and 7Be neutrinos in a LXe detector, as given in Ref. [18].
When the xenon atomic effects are considered (adopted in this paper), the rate in the region of interest (ROI) is suppressed by a few percent compared to the free electron scenario.
The PandaX-4T detector is located in the B2 hall of the China Jinping Underground Laboratory [19, 20].
The sensitive target of PandaX-4T contains 3.69 tonne of liquid xenon in a cylindrical dual-phase xenon time projection chamber (TPC) with 118.5 cm in diameter and 116.8 cm in height [21].
@@
This analysis selects different data samples compared to existing analyses of the commissioning data release [10, 13, 22].
The energy ROI is from 24 keV to 144 keV, with more than 60% of the solar neutrino events included.
The lower boundary is chosen to be above the dark matter particle search region and the upper boundary is to avoid the prominent 163.9 keV peak from 131mXe produced in the neutron calibration runs.
@@
Therefore, we can recover approximately 9.5% of exposure previously excluded in the dark matter analysis due to elevated detector noise [10].
On the contrary, 8.4 days of data following a neutron calibration run is removed because of a high concentration of activated 133Xe and 125I.
In the end, a total data set of 86.5 days is used for this analysis.
The PMT gains are calibrated with a newly implemented “rolling gain” approach, which fits single photoelectron (SPE) spectra in individual PMTs run by run, adopted already in recent Ref. [21].
@@
To avoid biases in our data selections, selection cuts are first determined with LED-gain calibration data, validated with about 13.5 days of rolling gain physics data, and finally applied to the entire data set.
The quality cut variables are inherited from the dark matter analysis [10], but cut parameters are modified
to suit the differing energy window.
@@
The scattering of solar neutrinos on electrons is primarily single-site (SS).
@@
The SS efficiency is , calculated using the 220Rn calibration and consistent with simulation.
@@
The reconstruction uniformity is confirmed with the diffusive 83mKr calibration source injected into the TPC [23].
The mono-energetic peak of 83mKr is also used to generate an energy response map. The corresponding correction procedure follows Ref. [10].
@@
The FV selection is the same as in the dark matter analysis geometrically, and a total of tonne of LXe in the center of the TPC is used.
@@
The primary challenge in this analysis lies in the accurate rate estimation of various background components with smooth energy spectra that resemble the shape of the solar signals.
@@
One major background in the ROI arises from the progeny of the internal 222Rn – the decay of 214Pb to the ground state of 214Bi [21].
The ratio of the events in the ROI to the full SS energy spectrum of 214Pb decay is determined by a dedicated 222Rn injection calibration run, with 214Pb SS events accumulated over approximately 11 days.
@@
The corresponding number of 214Pb events in the ROI is , in which the uncertainty combines both the fitted uncertainty of the overall 214Pb decay rate in the physics data [21] and the uncertainty of the event ratio within our ROI. Also along the decay chain of 222Rn, the decay of 214Bi can be identified effectively using the 214Bi–214Po coincidence, as the half-life of 214Po is 164 s.
In our analysis, when a 214Bi-like event is found, we reject the event if an event is found within the following 5 ms window. The residual 214Bi is no longer included in the spectrum fit. The loss of live time due to random coincidence is negligible.
The expected contribution from the decay of 212Pb in the 220Rn decay chain follows the approach in Ref. [21].
The ratio between 212Pb and the subsequent 212Po events is determined using a 220Rn injection calibration run. Then using the 212Po events identified in the physics data,
212Pb events in the ROI is estimated to be , where the uncertainty is dominated by the variation of 212Pb/212Po ratio during the 220Rn injection run.
The concentration of 85Kr is estimated to be 0.52 0.27 parts per trillion [21], leading to 489 254 events in the ROI.
The levels of radioactivity from PMTs and detector vessels have been determined from the wide energy spectrum fit in Ref. [22], resulting in events. The expected background from 136Xe is in the ROI, given by the half-life measured by the PandaX-4T experiment with an exposure of 15.5 kgyear of 136Xe isotope [22].
The contributions of 133Xe, 127Xe, 125I, and 124Xe are free and fitted in the final spectral analysis, as they cannot be constrained by our data outside the ROI.
The energy spectrum of 133Xe, activated during a neutron calibration run, has a distinctive rising slope starting at 81 keV.
Cosmogenic 127Xe, introduced when some xenon exposed at the Earth’s surface was filled into the detector, contributes to the background with K-shell electron capture around 33.2 keV.
Neutron-activated 125Xe decays quickly to
the relatively long-lived 125I with a half-life of 59.4 days.
125I electron capture can release a total energy of 67.3 keV (K-shell), 40.4 keV (L-shell), and 36.5 keV (M-shell) [27]. The reduction rate of 125I observed in the TPC is significantly more rapid than the natural lifetime of 125I, likely due to the removal by the xenon circulation and purification system.
Two-neutrino double-electron capture of 124Xe (abundance = 0.095%, Q = 2857 keV) deposits energy within our ROI at 64.3 keV (KK-shell) and 32.3 - 37.3 keV (KL, KM, and KN-shells) [28, 27]. Therefore, 125I (124Xe) events in the data are modeled as three (four) Gaussian peaks with relative areas fixed by the capture fractions and widths fixed by the resolution function (Fig. 2), but the overall normalization as a free fit parameter.
Table 1: Expected and fitted background and signal events in the ROI. Note that the fitted uncertainties contains both the statistical and some systematic components. See text for details.
@@
The contributions from 214Pb, 212Pb, 85Kr, detector materials, and 136Xe are constrained using the uncertainties described earlier and listed in Table 1.
The normalizations of 133Xe, 127Xe, 125I, and 124Xe are kept free in the fit.
The sum of solar and 7Be neutrino signals are combined into a single fit component.
The result of the spectrum fit is shown in Fig. 4.
The number of solar + 7Be neutrino signals is 231 257 events.
@@
The best-fit background contributions from detector materials, 85Kr, 214Pb, and 212Pb are very close to the input nominal values, indicating that the spectrum fit does not try to drive these background levels up and down, since their shape characters are similar.
The total uncertainty of the + 7Be neutrino signals is broken down into three terms (Table 2), the statistical uncertainty (), the systematic uncertainty incorporated in the likelihood fit with nuisance parameters (), and additional systematic uncertainty evaluated by hand (). We shall discuss them in turn below.
The statistical uncertainty is determined to be 113 events, by re-fitting the data with all constrained parameters fixed to their best-fit values without the penalty terms.
The overall size of is then evaluated as to be 231 events, representing the residual component of the fit uncertainty. To quantify the contribution to due to a given component (assuming that it is uncorrelated with the rest), we use a similar approach by refitting the data by only “tuning on” the nuisance parameter of this component, and all other nuisance parameters fixed to their best-fit values. The resulting fit uncertainty is then subtracted by in quadrature. Individual contributions thus evaluated are also summarized in Table 2. Among all constrained backgrounds, 85Kr, 214Pb and 212Pb dominate the contribution to with an uncertainty of 202, 87, and 69 events, respectively.
The uncertainties of data selection, including the FV, quality cuts, and the SS fraction, affect the signal number proportionally with an estimated uncertainty of 29 events.
Additional systematic uncertainties are evaluated manually.
The uncertainty from the reconstructed energy scale is determined by comparing the baseline fit result with new fit results, obtained by shifting the data energy spectrum by keV.
The larger difference led by the shifts is 142, which is used as the uncertainty.
To evaluate the contribution due to the energy resolution uncertainty, we perform the fit by imposing different background spectra with different energy resolution.
A difference in the energy resolution function propagates into an uncertainty of 19 events.
The fit range uncertainty is evaluated to be 29 events by varying the ROI to 25 – 143 keV and 23 – 145 keV.
The uncertainty introduced by the 214Pb spectrum shape estimated to be 84 events, by comparing the theoretical spectrum in the baseline fit to the shape directly measured in the calibration data.
Similarly, the spectral shape differences between the Geant4 [30] prediction and a recent theoretical calculation of 212Pb and 85Kr [26] are also considered, and the systematic uncertainties are found to be 18 and 5, respectively.
We also fit the data by taking the half-life of 136Xe from EXO-200 [32], the most precise result so far, as the input value.
This results in a difference of 16 events in comparison to our baseline fit, which is also taken as a component of systematic uncertainty. In total, is 170 events.
Table 2: Summary of contributions to the + 7Be neutrino signal uncertainties. The contribution of each constrained parameter is extracted by turning on/off the corresponding penalty term in the fitter, assuming no correlation with other constraints. As a result, the quadrature combination of those rows is slightly different from the subtotal obtained directly by the fitter with all correlations properly built in.
@@
29
214Pb spectrum
84
212Pb spectrum
18
85Kr spectrum
5
136Xe half-life
16
@@
287
Figure 4: (Color online) Result of the spectrum fit and the corresponding residual plot in the bottom panel. The solar +7Be neutrino is shown as the red line.
@@
Therefore, the final detected number of + 7Be neutrinos is events from our analysis.
The corresponding solar neutrino flux is , using the expected ratio of the to 7Be fluxes from SSM and their contributions to the ROI.
@@
Based on this, we obtain a neutrino flux upper limit of at 90% C.L..
This analysis represents the first result to directly measure the solar neutrinos at an electron recoil energy below 150 keV. It demonstrates the potential of multi-tonne-scale LXe detectors for solar neutrino studies in a completely new energy window.
@@
Additional efforts, such as replacing TPC materials and circulation pumps, are being implemented for background reduction.
Additional 222Rn calibration data and improved detector response with upgraded PMT readout circuit boards are expected to significantly reduce systematic uncertainties.
Assuming a 222Rn level of 3.5 Bq/kg, a krypton-to-xenon mole concentration of 0.25 ppt, and no short-lived xenon isotopes being activated, and imposing a 5% constraint on all backgrounds, PandaX-4T can measure the solar neutrinos with an uncertainty of 30% with 6 tonneyear exposure.
Figure 5: The solar neutrino flux measured by PandaX-4T (red), in comparison to those from gallium experiments [34] and Borexino [9].
@@
The PandaX-4T projection with 6 tonneyear exposure is indicated at the bottom (centered at the SSM).
This project is supported in part by the grants from National Science Foundation of China (Nos. 12090060, 12090063, 12105052, 12005131, 11905128, 11925502), the Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100),
and National Science Foundation of Sichuan Province (No. 2024NSFSC1371).
@@
Anselmann et al. [1992]P. Anselmann et al. (GALLEX), Solar neutrinos observed by GALLEX at Gran
Sasso., Phys. Lett. B 285, 376 (1992).